Solve for x
(x-1)/(x-5) = (2x+1)/(x-2)

Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

Solve for x
(x-1)/(x-5) = (2x+1)/(x-2)
-----
Cross-multiply to get:
(x-1)(x-2) = (x-5)(2x+1)
---
x^2 - 3x + 2 = 2x^2 -9x - 5
---------
x^2-6x -7 = 0
----
Factor:
(x-7)(x+1) = 0
x = 7 or x = -1
====================
Cheers,
Stan H.

Answer by lwsshak3(11628)

Solve for x
(x-1)/(x-5) = (2x+1)/(x-2)
LCD: (x-5)(x-2)
(x-1)(x-2)=(2x+1)(x-5)
you can also cross-multiply to get the same result
x^2-3x+2=2x^2-9x-5
x^2-6x-7=0
(x-7)(x +1)=0
x-7=0
x=7
or
x+1=0
x=-1
..
check:
for x=7x
(x-1)/(x-5) = (2x+1)/(x-2)
7-1)/(7-5) = (2*7+1)/7-2)
6/2=15/5
3=3
..
for x=-1
(-1-1)/(-1-5) = (2*-1+1)/(-1-2)
-2/-6=-1/-3
1/3=1/3